Use implicit differentiation to find on sin (2y?) - 5x = 5 ey Enter your answer in the answer box.
Answer #1

- Given that, Sin (24) -5x = 5ey Differentiating o with respect to a implicitly, we gete sing 2.yu) - 5x] = (524)ਜੋ ਰੰਤ [h (21) - ਰੰਕ (54) = ਰੰਕ (54)>cos2y) da (290)-5da. (a) = 5 taley) > c05/2ya).[2 400 - 5. (1) =5[evitaly cos(299[2424. ( 1177-5 = sebe * cos(2y9 [ og ] -S = 4 - (ਸ) ਲਾਹ R 6z dy 15 Ay cos (24) - 5ey